OpenStudy (123ab456c):

What is the value of 3√4096? A. 512 B. 64 C. 16 D. 8

OpenStudy (anonymous):

calculator exercise http://www.wolframalpha.com/input/?i=cuberoot%284096%29

OpenStudy (unklerhaukus):

4096 = 4000 + 96 = 20^2 + (100-4) = 20^2 + (10-2)(10+2)